40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  QIL elog  Not logged in ELOG logo
Message ID: 714     Entry time: Mon Apr 5 14:46:05 2010
Author: Alastair 
Type: Laser 
Category: GYRO 
Subject: CCW locking 

I've realigned the CCW mode and it is now locking reasonably.  The cavity had become misaligned and I had to adjust it to get it back on resonance.  I then swept the laser and maximised the peaks in transmission.  Next I put the PD back in for reflection locking.  I had to replace the mirror mounts with lens mounts because of the lack of room.  I adjusted the PD position to get maximum DC power, and then plugged it back into the PDH box.  We are getting dips that are approx 50% in voltage.  I couldn't get a decent error signal after the mixer until I reduced the sideband frequency (it's at 9MHz just now) but at the lower frequency we do get an error signal.  It's now locking reasonably but needs more work on the servo settings to get back to where we were before.

ELOG V3.1.3-